Find enenrgy of photon whose wavelength is 3.8 x 10-7m
1
Expert's answer
2020-12-15T11:32:41-0500

"E=\frac{h\times c}{\lambda}=\frac{6.6\times10^{-34}\times 3\times10^8}{3.8\times 10^{-7}}=5.21\times 10^{-19}" J